# Generative modeling of conditional probability distributions on the level-sets of collective variables

### v2

@@ Abstract
Given a probability distribution $\mu$ in $\mathbb{R}^{d}$ represented by data, we study in this paper the generative modeling of its conditional probability distributions on the level-sets of a collective variable $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$. We propose a general and efficient learning approach that is able to learn generative models on different level-sets of $\xi$ simultaneously. To improve the learning quality on level-sets in low-probability regions, we also propose a strategy for data enrichment by utilizing data from enhanced sampling techniques. We demonstrate the effectiveness of our proposed learning approach through concrete numerical examples. The proposed approach is potentially useful for the generative modeling of molecular systems in biophysics, for instance.
@@ 1 Introduction
High-dimensional stochastic processes are ubiquitous in broad research areas such as biophysics, materials science, and climate modeling. Understanding the complex dynamical behavior of such processes often requires projecting the system down to a (much) lower-dimensional space via a few features, i.e. collective variables (CVs) or reaction coordinates, of the system. In molecular dynamics, for instance, various enhanced sampling techniques have been developed to overcome the sampling difficulty due to the system’s metastability and to construct the free energy surface in the space of CVs [15]. A number of analytical and computational approaches have also been proposed for constructing simpler, lower-dimensional surrogate models associated to a chosen set of CVs of the original high-dimensional system. One notable approach is the effective dynamics constructed using conditional expectations, where the coefficients (i.e. drift and diffusion) in the surrogate model are computed by taking averages with respect to the so-called conditional probability distribution on the corresponding level-set of the CVs [21]. This approach has attracted considerable research interest in the mathematical community [50, 51, 22, 29, 23, 9, 52], contributing to a deeper understanding of the quality of the effective dynamics in approximating the original high-dimensional processes.
@@
A related research topic in molecular dynamics is sampling on the level-sets of CVs [5, 13, 26], which is useful in both of the aforementioned tasks, i.e. free energy computation and computing the coefficients in the effective dynamics. Metropolis-Hastings Monte Carlo methods have been developed for unbiased sampling on submanifolds defined by the level-sets of CVs [48, 27, 28]. Given a (Boltzmann–Gibbs) distribution in the full space, sampling methods for its conditional distribution on the level-sets of CVs have been developed by constructing ergodic diffusion processes on the level-sets [49, 42]. More broadly, methods for sampling under constraints have also been studied in computational statistics [46, 35].
@@
In recent years, generative modeling approaches, in particular diffusion models [43, 17, 44, 45] and flow-matching models [31, 32, 1], have achieved significant success in a wide range of real-world applications. Motivated by these achievements, there has been growing research interest in extending generative modeling approaches for Euclidean data to manifold data [36, 41, 2, 8, 18, 19, 4, 34, 33]. However, to the best of our knowledge, most of these existing approaches rely on substantial geometric information such as geodesics [8], heat kernel, or eigenfunctions [4], which consequently limits their applicability to manifolds where such information is explicitly known. Only the methods proposed in [34, 33] can be applied to manifolds that are implicitly defined as a level-set of a certain function. All these approaches are developed for generative modeling on a single and fixed manifold.
@@
In this paper, we study the generative modeling of conditional probability distributions on the level-sets of CVs and propose a versatile framework for this task. Our approach builds upon the flow-matching model [31, 32, 1], in which a neural ordinary differential equation (ODE) is learned to transport samples from a base Gaussian distribution to the target conditional distribution on each level-set of a CV map $\xi$. Our method does not require explicit geometric information of the underlying manifolds and therefore is broadly applicable to a wide range of manifolds implicitly defined by CVs. Moreover, by varying the target CV value, the method can generate samples across different level-sets of $\xi$, rather than being restricted to a single level-set. To further improve learning in regions of low probability density, we introduce a data enrichment strategy that leverages data obtained from enhanced sampling techniques such as adaptive biasing force (ABF) simulations. The approach is validated on a hierarchy of systems of increasing complexity, from synthetic low-dimensional examples to high-dimensional molecular datasets. The main contributions of this paper are summarized below:
@@
    comprehensive numerical validations that demonstrate the flexibility, robustness, and practical applicability of the method to molecular systems.
@@ 2 Background and notation
The aim of this section is to introduce the probability distributions of interest in this work, in particular the conditional distribution on the level-sets of a given smooth map, which will play a central role in the following.
Consider a target distribution $\mu$ over $\mathbb{R}^{d}$ with a positive probability density $\rho$ with respect to the Lebesgue measure, namely $\mu(dx)=\rho(x)\,dx$. Assume that a $C^{2}$-smooth map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is given, where $1\leq k<d$. We are particularly interested in applications in statistical physics, where $\mu$ arises as the invariant distribution of certain ergodic diffusion processes $x_{t}$ in $\mathbb{R}^{d}$ and $\xi$ corresponds to the CV of the system. In this case, we often refer to the Boltzmann–Gibbs distribution, whose density with respect to the Lebesgue measure is
@@
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that grows sufficiently fast to infinity as $|x|\rightarrow+\infty$, $\beta$ is a positive constant whose inverse is proportional to the system’s temperature, and $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,dx$ is the normalizing constant. We introduce a few quantities related to the measure $\mu$ and the map $\xi$. Further discussions can be found in [25, Section 3.2.1] and [50].
Given $z\in\mathbb{R}^{k}$, we define $\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}$ as the level-set of the map $\xi$. Assume that $\Sigma_{z}$ is non-empty and that the Jacobian of $\xi$ at $x$, denoted by $\nabla\xi(x)\in\mathbb{R}^{d\times k}$, has full rank $k$ at each $x\in\Sigma_{z}$. Under this assumption, the regular value theorem implies that $\Sigma_{z}$ is a (dk)(d-k)-dimensional submanifold of $\mathbb{R}^{d}$; see [16, Chapter 1, Theorem 3.2]. Let $\widetilde{\mu}$ denote the pushforward measure of $\mu$ by the map $\xi$; that is, $\widetilde{\mu}$ is the probability measure of the random variable $\xi(X)\in\mathbb{R}^{k}$ when $X\sim\mu$. Let $\mu_{z}$ denote the conditional probability measure on $\Sigma_{z}$. The law of total expectation implies
$\mathbb{E}_{\mu}(f(X))=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu}\big(f(X)\mid\xi(X)=z\big)\Big]=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu_{z}}\big(f(X)\big)\Big]\,$, (2)
for a test function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, the co-area formula implies that (see [10, Theorem 3.11] and [25, Lemma 3.2])
@@
is a normalizing factor in (3). For processes whose invariant distribution is the Boltzmann–Gibbs distribution $\mu(dx)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}dx$, the free energy associated with the map $\xi$ is defined as
@@ 3 Learning approach
In this section, we present our methodology for learning generative models on the level-sets of a smooth map $\xi$. We first describe the standard approach based on flow-matching for learning the conditional distributions $\mu_{z}$ corresponding to different values $z\in\mathbb{R}^{k}$. Then, we discuss an enhanced learning strategy that leverages biased trajectory data to improve the quality of the learning in regions of low probability under the pushforward distribution $\widetilde{\mu}$. Finally, we explain how such biased trajectory data can be generated in practice using adaptive biasing techniques.
@@ 3.1 Standard learning approach
Our goal is to sample the conditional probability measure $\mu_{z}$ on the level-sets of $\xi$ for different values $z\in\mathbb{R}^{k}$. For this purpose, we adopt the flow-matching generative modeling framework [32, 31, 1], which employs ordinary differential equations (ODEs) in $\mathbb{R}^{d}$
@@
$\mathcal{L}_{\mathrm{FM},\nu}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\nu}\left\|v\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,.$ (9)
It is known that the minimizer of (9) leads to an ODE (8) that transforms the prior distribution $\nu_{\mathrm{prior}}$ to the target $\nu$.
Adapting this framework to our setting, we propose to learn the ODEs
@@
parametrized by a variable $z\in\mathbb{R}^{k}$, such that, for a fixed $z\in\mathbb{R}^{k}$, the system (10) transforms the prior distribution from $t=0$ to the conditional measure $\mu_{z}$ at $t=1$. To achieve this goal, we consider the following objective for learning the vector field in (10):
@@
where $t_{1},t_{2},\dots,t_{N}$ are drawn independently from $U([0,1])$, $X_{0}^{(1)},X_{0}^{(1)},\dots,X_{0}^{(N)}$ are drawn independently from $\nu_{\mathrm{prior}}$, and $X^{(1)},X^{(2)},\dots,X^{(N)}$ are the data points.
@@ 3.2 Enhanced learning approach with biased trajectory data
We consider the case where the target distribution $\mu$ is the Boltzmann–Gibbs distribution in (1), which is the invariant distribution of a diffusion process $x_{t}$ in $\mathbb{R}^{d}$. For simplicity, we assume that $x_{t}$ is governed by the overdamped Langevin dynamics
@@
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that increases to infinity sufficiently fast as $|x|\rightarrow+\infty$ and $w_{t}$ is a standard Brownian motion in $\mathbb{R}^{d}$. In this case, it is known that the Boltzmann–Gibbs distribution $\mu$ in (1) is invariant under the process $x_{t}$ in (14) and that $x_{t}$ is ergodic with respect to $\mu$. Thanks to the ergodicity of the process, the empirical objective can be evaluated from trajectory data $X^{(1)},X^{(2)},\dots,X^{(N)}$ obtained by sampling (14) using a numerical scheme at time $t=h,2h,\dots,Nh$, where $h$ denotes the step-size and $N$ is the total number of sampling steps.
We are particularly interested in the case where the potential $V$ has multiple local minima and $\beta$ is (relatively) large. In this case, the probability density of the invariant distribution $\mu$ is highly non-uniform in space, since the probability that $x_{t}$ is within the vicinity of one of the local minima (metastable regions) is much larger than the probability that $x_{t}$ is elsewhere. Likewise, for a proper CV map $\xi$, the pushforward distribution $\widetilde{\mu}$ of $\mu$ defined in (4) is highly non-uniform in $\mathbb{R}^{k}$ as well. In this case, the generative model that is learned by minimizing the objective (11) (or its empirical counterpart (13)) using training data sampled from (14) may be inferior for values $z$ that belong to low-density regions under $\widetilde{\mu}$, due to the relatively small contribution of the loss in such regions to the aggregated objective (11) (see (12)).
In order to overcome the aforementioned issue, we utilize the observation (as can be seen from the derivation (12)) that different probability distributions of $X_{1}$ can be chosen in (11), as long as the corresponding conditional probability distributions coincide with the conditional probability distributions $\mu_{z}$ of $\mu$. In particular, let us consider the reweighted (RW) probability distribution defined by
@@
where $A:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\bar{Z}=Z^{-1}\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx$ is a normalizing factor. Then, for each $z\in\mathbb{R}^{k}$, it is straightforward to verify that the conditional probability distribution of $\mu_{\mathrm{RW}}$ on the level-set $\Sigma_{z}$ is $\mu_{z}$ given in (3). Moreover, using (5) and (6), we can obtain that the pushforward of $\mu_{\mathrm{RW}}$ by the map $\xi$ satisfies
@@
where $F$ is the free energy (6) associated to the map $\xi$. With the choice of the probability distribution $\mu_{\mathrm{RW}}$ in (15), the objective (11) becomes
@@
and by analogy with (18) we can derive that
@@
$\frac{\partial F}{\partial z_{i}}(z)=\mathbb{E}_{\mu_{z}}\left[\sum_{l=1}^{d}\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\frac{\partial V}{\partial x_{l}}-\beta^{-1}\sum_{l=1}^{d}\frac{\partial}{\partial x_{l}}\bigg(\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\bigg)\right]\,,\quad i=1,2,\dots,k\,$, (21)
where $(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}$ denotes the $ij$-component of the inverse of the matrix $\nabla\xi^{\top}\nabla\xi\in\mathbb{R}^{k\times k}$. In practice, a grid in $\mathbb{R}^{k}$ is defined and the mean value (21) is estimated for each cell of the grid. In each simulation step, the estimation is updated for the cell which the current state $\xi(x_{t})$ falls into. When the cell has been visited for more than a certain (pre-defined) number of times, the current value of the estimation is used as the biasing term $f_{t}(\xi(x_{t}))$ in (20). As $t\rightarrow+\infty$, we can expect that the empirical estimations converge to the true means in (21) and, accordingly, the process (20) converges to its equilibrium counterpart (19) with $A=F$.

### v3

@@ Abstract
Given a probability distribution $\mu$ in $\mathbb{R}^{d}$ represented by data, we study in this paper the generative modeling of the corresponding conditional probability distributions on the level-sets of a collective variable $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$. We propose a general and efficient learning approach that can learn generative models on different level-sets of $\xi$ simultaneously. To improve the learning quality on level-sets in low-probability regions, we also propose a data enrichment strategy by utilizing data from enhanced sampling techniques. We demonstrate the effectiveness of our proposed learning approach through concrete numerical examples. The proposed approach is potentially useful for the generative modeling of molecular systems in biophysics.
@@ 1 Introduction
High-dimensional stochastic processes are ubiquitous in broad research areas such as biophysics, materials science, and climate modeling. Understanding the complex dynamical behavior of such processes often requires projecting the system down to a (much) lower-dimensional space defined by a few features, i.e. collective variables (CVs) or reaction coordinates, of the system. In molecular dynamics, for instance, various enhanced sampling techniques have been developed to increase the sampling efficiency and to construct the free energy surface in the space of CVs [15]. A number of analytical and computational approaches have also been proposed for constructing simpler, lower-dimensional surrogate models associated to a chosen set of CVs of the original high-dimensional system. A notable approach is the effective dynamics constructed using conditional expectations, where the coefficients (i.e. drift and diffusion) in the surrogate model are computed by taking averages with respect to the so-called conditional probability distribution on the corresponding level-set of the CVs [21]. This approach has attracted considerable research interest in the mathematical community [50, 51, 22, 29, 23, 9, 52], contributing to a deep understanding on the quality of the effective dynamics in approximating the original high-dimensional processes.
@@
A common task in both of the aforementioned research directions, i.e. free energy calculation and the computation of coefficients in the effective dynamics, is sampling on the level-sets of CVs [5, 13, 26]. Metropolis-Hastings Monte Carlo methods have been developed for unbiased sampling on submanifolds defined by the level-sets of CVs [48, 27, 28]. Given a (Boltzmann–Gibbs) distribution in the full space, sampling methods for the corresponding conditional distribution on the level-sets of CVs have been developed by constructing ergodic diffusion processes on the level-sets [49, 42]. More broadly, methods for sampling under constraints have also been studied in computational statistics [46, 35].
@@
In recent years, generative modeling approaches, in particular diffusion models [43, 17, 44, 45] and flow-matching models [31, 32, 1], have achieved significant success in a wide range of real-world applications. Motivated by these achievements, there has been growing research interest in extending generative modeling approaches for Euclidean data to manifold data [36, 41, 2, 8, 18, 19, 4, 34, 33]. Most of these existing approaches rely on substantial geometric information such as geodesics [8], heat kernel, or eigenfunctions [4]. As a result, their applicability is limited to manifolds where such information is explicitly known. Only the methods proposed in [34, 33] can be applied to manifolds that are implicitly defined as a level-set of a certain function. However, to the best of our knowledge, all these approaches are developed for generative modeling on a single and fixed manifold.
@@
In this paper, we propose an effective generative modeling approach for the conditional probability distributions on the level-sets of CVs. Our approach builds upon the flow-matching model [31, 32, 1], in which a neural ordinary differential equation (ODE) is learned to transport a prior (Gaussian) distribution to the target conditional distributions on the level-sets of a CV map $\xi$. Our method does not require explicit geometric information of the underlying manifolds and is broadly applicable to a wide range of manifolds that are implicitly defined by CVs. By varying the target CV value, the method can generate samples across different level-sets of $\xi$, rather than being restricted to a single level-set. To further improve the learning in regions of low probability density, we introduce a data enrichment strategy that leverages data obtained from enhanced sampling techniques such as adaptive biasing force (ABF) simulations. The approach is validated on a hierarchy of systems of increasing complexity, from synthetic low-dimensional examples to high-dimensional molecular datasets. The main contributions of this paper are summarized below:
@@
    Comprehensive numerical validations that demonstrate the flexibility, robustness, and practical applicability of the method.
@@ 2 Background and notation
The aim of this section is to introduce the probability distributions of interest in this work, in particular the conditional distributions on the level-sets of a given smooth map.
Consider a distribution $\mu$ over $\mathbb{R}^{d}$ with a positive probability density $\rho$ with respect to the Lebesgue measure, namely $\mu(dx)=\rho(x)\,dx$. Assume that a $C^{2}$-smooth map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is given, where $1\leq k<d$. We are particularly interested in applications in statistical physics, where $\mu$ arises as the invariant distribution of certain ergodic diffusion process $x_{t}$ in $\mathbb{R}^{d}$ and $\xi$ corresponds to the CV of the system. In this case, we often refer to the Boltzmann–Gibbs distribution, whose density with respect to the Lebesgue measure is
@@
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that grows sufficiently fast to infinity as $|x|\rightarrow+\infty$, $\beta$ is a positive constant whose inverse is proportional to the system’s temperature, and $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,dx$ is the normalizing constant that we assume to be finite. Below we introduce a few quantities related to the measure $\mu$ and the map $\xi$. Further discussions can be found in [25, Section 3.2.1] and [50].
Given $z\in\mathbb{R}^{k}$, the level-set of the map $\xi$ corresponding to $z$ is defined as
$\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}.$
Assume that $\Sigma_{z}$ is non-empty and that $\nabla\xi(x)\in\mathbb{R}^{d\times k}$, the Jacobian of $\xi$, has full rank $k$ at each $x\in\Sigma_{z}$. Under these assumptions, the regular value theorem asserts that $\Sigma_{z}$ is a (dk)(d-k)-dimensional submanifold of $\mathbb{R}^{d}$; see [16, Chapter 1, Theorem 3.2]. Let $\widetilde{\mu}$ denote the pushforward measure of $\mu$ by the map $\xi$; that is, $\widetilde{\mu}$ is the probability distribution of the random variable $\xi(X)\in\mathbb{R}^{k}$ when $X\sim\mu$. Let $\mu_{z}$ denote the conditional probability distribution on $\Sigma_{z}$, i.e. the probability distribution of $X$ conditioned on the event that $\xi(X)=z$. The law of total expectation implies
$\mathbb{E}_{X\sim\mu}(f(X))=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu}\Big(f(X)\,\Big|\,\xi(X)=z\Big)\Big]=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu_{z}}\big(f(X)\big)\Big]\,$, (2)
for test functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, the co-area formula implies that (see [10, Theorem 3.11] and [25, Lemma 3.2])
@@
is a normalizing factor in (3). For processes whose invariant distribution is the Boltzmann–Gibbs distribution $\mu(dx)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}dx$, the free energy associated to the map $\xi$ is defined as
@@ 3 Learning approach
In this section, we present our methodology for learning generative models on the level-sets of a smooth map $\xi$. We first describe the standard approach based on flow-matching for learning the conditional distributions $\mu_{z}$ corresponding to different values $z\in\mathbb{R}^{k}$. Then, we propose an enhanced learning strategy that leverages biased trajectory data to improve the quality of the learning in regions of low probability under the pushforward distribution $\widetilde{\mu}$. Finally, we explain how such biased trajectory data can be generated in practice using adaptive biasing techniques.
@@ 3.1 Standard learning approach
Our goal is to sample the conditional probability distributions $\mu_{z}$ for different values $z\in\mathbb{R}^{k}$. For this purpose, we adopt the flow-matching generative modeling framework [32, 31, 1], which employs the ordinary differential equation (ODE) in $\mathbb{R}^{d}$
@@
$\mathcal{L}_{\mathrm{FM},\nu}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\nu}\left\|v\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,$, (9)
where $U([0,1])$ denotes the uniform distribution on $[0,1]$. It is known that the minimizer of (9) leads to an ODE (8) that transforms the prior distribution $\nu_{\mathrm{prior}}$ to the target $\nu$.
Adapting this framework to our setting, we propose to learn the ODE
@@
parametrized by a variable $z\in\mathbb{R}^{k}$, such that, for a fixed $z\in\mathbb{R}^{k}$, the system (10) transforms the prior distribution from $t=0$ to the conditional distribution $\mu_{z}$ at $t=1$. To achieve this goal, we consider the following objective for learning the vector field in (10):
@@
where $t_{1},t_{2},\dots,t_{N}$ are drawn independently from $U([0,1])$, $X_{0}^{(1)},X_{0}^{(1)},\dots,X_{0}^{(N)}$ are drawn independently from the prior $\nu_{\mathrm{prior}}$, and $X^{(1)},X^{(2)},\dots,X^{(N)}$ are the data points.
@@ 3.2 Enhanced learning approach with biased trajectory data
We consider the case where the target distribution $\mu$ is the Boltzmann–Gibbs distribution with a density in (1). Let us introduce the overdamped Langevin dynamics
@@
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that increases to infinity sufficiently fast as $|x|\rightarrow+\infty$ and $w_{t}$ is a standard Brownian motion in $\mathbb{R}^{d}$. It is known that $\mu$ is invariant under the process $x_{t}$ in (14) and that $x_{t}$ is ergodic with respect to $\mu$. Thanks to the ergodicity of the process, the empirical objective (13) can be evaluated using trajectory data $X^{(1)},X^{(2)},\dots,X^{(N)}$ obtained by sampling (14) with a numerical scheme at time $t=h,2h,\dots,Nh$, where $h$ denotes the step-size and $N$ is the total number of sampling steps.
We are particularly interested in the case where the potential $V$ has multiple local minima and $\beta$ is relatively large (i.e. the temperature is low). In this case, the probability density of the invariant distribution $\mu$ is highly non-uniform in space, since the probability that $x_{t}$ is within the vicinity of one of the local minima (metastable regions) is much larger than the probability that $x_{t}$ is found elsewhere. Likewise, for a proper CV map $\xi$, the pushforward distribution $\widetilde{\mu}$ defined in (4) is highly non-uniform in $\mathbb{R}^{k}$ as well. The generative model that is learned by minimizing the objective (11) (or its empirical counterpart (13)) using training data sampled from (14) may be inferior in the low-density regions under $\widetilde{\mu}$, because of the relatively small contribution of the loss in such regions to the aggregated objective (11) (see (12)).
In order to overcome the aforementioned issue, we utilize the observation (as can be seen from the derivation (12)) that different probability distributions of $X_{1}$ can be chosen in (11), as long as the corresponding conditional probability distributions coincide with the conditional probability distributions $\mu_{z}$. In particular, let us consider the reweighted (RW) probability distribution
@@
where $A:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\bar{Z}=Z^{-1}\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx$ is a normalizing factor. Then, for each $z\in\mathbb{R}^{k}$, it is straightforward to verify that the conditional probability distribution of $\mu_{\mathrm{RW}}$ on the level-set $\Sigma_{z}$ coincides with $\mu_{z}$ given in (3). Moreover, using (5) and (6), we can obtain that the pushforward of $\mu_{\mathrm{RW}}$ by the map $\xi$ satisfies
@@
where $F$ is the free energy (6) associated to the map $\xi$. With the choice of the probability distribution $\mu_{\mathrm{RW}}$ in (15), the objective (11) changes to
@@
and similar to (12) we can derive that
@@
$\frac{\partial F}{\partial z_{i}}(z)=\mathbb{E}_{\mu_{z}}\left[\sum_{l=1}^{d}\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\frac{\partial V}{\partial x_{l}}-\beta^{-1}\sum_{l=1}^{d}\frac{\partial}{\partial x_{l}}\bigg(\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\bigg)\right]\,,~i=1,2,\dots,k\,$, (21)
where $(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}$ denotes the $ij$-component of the inverse of the matrix $\nabla\xi^{\top}\nabla\xi\in\mathbb{R}^{k\times k}$. In practice, the mean value (21) is estimated on a grid in $\mathbb{R}^{k}$. In each simulation step, the estimation is updated for the cell where the current state $\xi(x_{t})$ belongs to. After the cell has been visited for a certain (pre-defined) number of times, the current value of the estimation in that cell is used as the biasing term $f_{t}(\xi(x_{t}))$ in (20). As $t\rightarrow+\infty$, the empirical estimations are expected to converge to the true means in (21) and, accordingly, the process (20) is expected to converge to its equilibrium counterpart (19) with $A=F$.